Help me find the discriminant for :
3x^2-14x+9

Answer by stanbon(48546)

find the discriminant for :
3x^2-14x+9
------
a=3 ; b=-14; c=9
discriminant = b^2-4ac
=14^2 - 4*3*9
=196 - 108
=88
========
Cheers,
Stan H.